OpenStudy (okdutchman7):

Please help. WILL MEDAL! Trigonometry @TheSmartOne @myininaya @nincompoop @jigglypuff314 @johnweldon1993 @jim_thompson5910 @mathstudent55 @zepdrix @Zarkon @calculusxy @iPwnBunnies

OpenStudy (okdutchman7):

\[3\cot \frac{ \theta }{ 2 } -\sqrt{3} = 0\]

OpenStudy (okdutchman7):

\[3\cot \frac{ \theta }{ 2 } = \sqrt{3}\]

OpenStudy (mathstudent55):

Divide both sides by 3.

OpenStudy (okdutchman7):

Do I just divide the argument of the cotangent function or the whole thing?

OpenStudy (okdutchman7):

\[\cot \frac{ \theta }{ 2 } = \frac{ \sqrt{3} }{ 3 }\]

OpenStudy (mathstudent55):

Divide both sides just by 3. This does not involve the argument of the cotangent. For example: \(\dfrac{4 \tan \frac{\theta}{2}}{4} = \tan \frac{\theta}{2}\)

OpenStudy (okdutchman7):

Is that ^ right then?

OpenStudy (mathstudent55):

\(3\cot \frac{ \theta }{ 2 } -\sqrt{3} = 0\) \(3\cot \frac{ \theta }{ 2 } = \sqrt{3} \) \(\cot \frac{ \theta }{ 2 } = \dfrac{\sqrt{3}}{3} \) This is correct. The division by 3 is not in the argument. It is only in the coefficient.

OpenStudy (okdutchman7):

What do I do next?

OpenStudy (mathstudent55):

The cotangent is the reciprocal of the tangent. \(\tan \frac{ \theta }{ 2 } = \dfrac{3}{\sqrt{3}} = \sqrt 3 \)

OpenStudy (okdutchman7):

tan is opposite/agacent

OpenStudy (mathstudent55):

Now we solve for \(\dfrac{\theta}{2} \). For what value of an angle is the tangent equal to \(\sqrt 3\) ?

OpenStudy (okdutchman7):

The angle is 60

OpenStudy (okdutchman7):

are there others?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Are you solving in an interval, or you need all values?

OpenStudy (okdutchman7):

I need all values between 0 and 360.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The tangent is positive in quadrants I and III.

OpenStudy (okdutchman7):

so 240

OpenStudy (okdutchman7):

Is the \[\tan \frac{ \theta }{ 2 }\] theta over 2 significant?

OpenStudy (mathstudent55):

We solved so far for \(\dfrac{\theta}{2} \), and we get that the tangent is \(\sqrt 3 \) at \(60^\circ + 180^\circ k\) We now have \(\dfrac{\theta}{2} = 60^\circ + 180^\circ k\) Now we solve for \(\theta\) by multiplying both sides by 2. \(\theta = 120^\circ + 360^\circ k\) Now we let k = 0, 1, 2, 3, ..., and we see which values will produce an answer in the interval [0, 360).

OpenStudy (okdutchman7):

so 120 is the only value between 0 and 360

OpenStudy (okdutchman7):

480 is too big

OpenStudy (mathstudent55):

\(\theta = 120^\circ + 360^\circ k\) For k = 0, \(\theta = 120^\circ + 360^\circ \times 0 = 120^\circ\) For k = 1, \(\theta = 120^\circ + 360^\circ \times 1 = 120^\circ + 360^\circ = 480^\circ\) which is not in the interval [0, 360). The only answer is \(\theta = 120^\circ\)

OpenStudy (mathstudent55):

Correct.

OpenStudy (okdutchman7):

Thanks!

OpenStudy (mathstudent55):

You're welcome.